Suppose that you are interested in the recent introduction of a non-native avian predator, a falcon, and how it has influenced the evolution of flight in one of its prey species, the song sparrow (Melospiza melodia). Just prior to introduction of the falcon, mean wing length of sparrows was 9.8 cm. Following one season of predation by falcons, mean wing length in the surviving sparrows was 9.1 cm.

1. If the mean wing length of offspring of these surviving sparrows is 9.5 cm, what is the narrow sense heritability for wing length (to the nearest thousandth)?

2.If the total phenotypic variance in wing length for the unselected parental population is VP = 3.2, then what is the additive genetic variance (VA) for this trait (to the nearest thousandth)?
Answer #1

1.Narrow sense heritability is the ratio of additive genetic variance to total phenotypic variance.

H2 =VA/ VP

H2 =9.5-9.1/9.8-9.1

=0.4/0.7

=0.571

=57.1%

2. H2 = VA/VP

VA = 0.57×3.2

=1.82